# Learning topological states from randomized measurements using variational tensor network tomography

### v3

@@ B.6 Main result: A concentration inequality for infidelity
For simplicity, we assume that $|\hat{\psi}^{(N)}\rangle$ and $|\phi\rangle$ are normalized elements of the equivalence classes $[\hat{\psi}^{(N)}]$ and $[\phi]$, respectively. Let $\tilde{g}$ be the Fubini-Study metric of the projective MPS manifold $\tilde{M}$ induced by the Fubini-Study metric of the projective Hilbert space via Eq. (41). On this manifold, $\gamma$ is the shortest connecting geodesic between $\gamma(0)=[\phi]$ and $\gamma(1)=[\hat{\psi}^{(N)}]$. By Proposition. 1, the estimator converges to the target state in the infinite-sample limit
$[\hat{\psi}^{(N)}]\overset{\mathrm{p.}}{\longrightarrow}[\phi],$ (64)
@@
[…]
@@
[…]
@@
The $\eta$-term in the denominator is a consequence of the fact that the estimator $[\hat{\psi}^{(N)}]$ is only asymptotically normal. The smaller one wants to make $\eta$, the larger $N_{0}$ becomes so that we have to move “closer” to the asymptotic regime. Asymptotically, i.e., in the limit $N\rightarrow\infty$, we can choose $\eta\rightarrow 0$ to arrive at our main result in Theorem 2
@@
Supplementary Figure 1: Infidelity scaling for interpolation between a random and GHZ $3$-qubit state. The numerical results illustrate the polynomial scaling of the infidelity with the number of samples $N$ across a family of states, parametrized by $x$. Here, $x$ ranges from 0 (darkest color, representing the GHZ state) to 1 (lightest color, representing the random state). Dashed lines indicate polynomial fits of the form $1-F=c\times N^{-\alpha}$, with values for $c$ and $\alpha$ provided in the legend.
@@ B.7 Discussions and interpretations
In this subsection, we discuss Theorem 2 and its connection to the numerical results presented in the main text. This theorem demonstrates that the MPS solution of maximum likelihood estimate (MLE) approaches the target state when the number of samples $N$ is sufficiently large, with the infidelity $1-F$ converging at a rate proportional to $1/\sqrt{N}$. This provides a theoretical foundation for understanding the scaling behavior observed in our numerical experiments.
First, we comment on how this theoretical result relates to our numerics in practice. In the derivation of our theorem, we have assumed that our variational space forms a projective MPS manifold. However, in our numerical simulations, we fix the bond dimension across sites, including variational states that are not always full rank, although they do include a projective MPS manifold as a subspace. In addition, several assumptions underpinning our result could be compromised: 1) The Hessian of the negative log-likelihood (NLL) loss function may not behave as expected. 2) The unitary ensemble is not tomographically complete in our variational space. Notably, while Theorem 1 confirms that our random-$XZ$ measurements are tomographically complete for real and pure states, our variational approach involves complex numbers for enhanced stability in optimization, leading to a space of complex pure states. To overcome this discrepancy, incorporating random-Pauli measurements, which are known to be tomographically complete, could provide a solution. Alternatively, our analysis of the complex projective MPS manifold could be adapted to the setting of a real projective MPS, and one can verify that the necessary conditions on the manifold structure are satisfied in the subspace of real states.
@@
[…]
@@
Random-$XZ$ classical shadows involve preparing some (unknown) state, randomly measuring some $n$-qubit Pauli string composed of only $X$ and $Z$ single-site Paulis, and then post-processing to predict properties of interest. We assume that the chosen Pauli is uniformly sampled from the set of all $n$-qubit strings composed of only $X$s and $Z$s. As this set of measurements is not informationally complete, random-$XZ$ classical shadows do not allow us to predict any arbitrary observable. However, we can predict some observables of interest, namely, those that contribute to our regularization step. In this appendix, we will discuss which observables can be estimated with this technique, discuss how to post-process the data to estimate observables, and how to calculate the sample complexity required for learning Pauli strings. We specifically discuss the sample complexity for learning Pauli strings because these are the observables that we estimate in our regularization step. In the first section of this appendix, we introduce these ideas pedagogically, with the rigorous backings for all our claims provided in the second section.
@@ C.1 Performing random-$XZ$ classical shadow tomography
$U=\otimes_{i=1}^{n}U^{\alpha_{i}}\hskip 2.84526pt\text{ where }\alpha_{i}\in\{X,Z\}.$ (87)
Here, $U^{\alpha_{i}}$ is the single-qubit unitary which rotates from the computational basis to the (different) basis of the Pauli operator $\sigma_{i}^{\alpha}$ acting on site $i$. Therefore, acting on the prepared state, the ensemble $\mathcal{U}_{XZ}$ allows one to randomly measure any $n$-qubit Pauli string composed of only $X$ and $Z$ single-site Paulis. Again, this set of measurements is not informationally complete, and as such, we cannot predict any arbitrary observable with random-$XZ$ classical shadows.
@@
[…]
@@
$=$ $\otimes_{i=1}^{n}\left(\mathop{\mathrm{Tr}}(U^{\alpha_{i}\dagger}\rvert b_{i}\rangle\!\langle b_{i}\lvert U^{\alpha_{i}}X)X+\mathop{\mathrm{Tr}}(U^{\alpha_{i}\dagger}\rvert b_{i}\rangle\!\langle b_{i}\lvert U^{\alpha_{i}}Z)Z+\frac{1}{2}\mathbb{I}\right)$
$=$ $\otimes_{i=1}^{n}\left(2U^{\alpha_{i}\dagger}\rvert b_{i}\rangle\!\langle b_{i}\lvert U^{\alpha_{i}}-\frac{\mathbb{I}}{2}\right).$
@@
Here, we represent the unitary applied in the $j$th random measurement and the outcome bit-string of that measurement as $U^{(j)}$ and $b^{(j)}$, respectively.
@@ C.2 Derivation of tools for random-$XZ$ classical shadows
[…]
@@
[…]
@@
[…]
@@ D.2 Proof of Theorem 1
[…]
@@
[…]
@@
[…]
$O=\otimes_{i=1}^{k}O_{i};\hskip 11.38109ptO_{i}=\begin{cases}\lvert+\rangle\!\langle+\rvert&\quad\text{if }i=k,\\ \left\rvert b^{(0)}_{i}\right\rangle\left\langle b^{(0)}_{i}\right\lvert&\quad\text{if }b^{(0)}_{i}=b^{(1)}_{i},\\ X&\quad\text{otherwise},\\ \end{cases}$ (102)
where $b^{(0)}_{i}$ is the $i$th bit of bit-string $b^{(0)}$, and likewise $b^{(1)}_{i}$ is the $i$th bit of $b^{(1)}$.
@@
[…]
@@ E.1 Maximum likelihood estimation: Randomized Kullback–Leibler divergence
More specifically, for our purposes, the dataset consists of bit-string measurements of a target state $\rvert\phi\rangle\!\langle\phi\lvert$ in the computational basis after applying a unitary $U$. In a randomized measurement scheme, such as the $XZ$-shadow measurements discussed in Appendix C, we first draw a unitary from an ensemble $\mathcal{U}$ with a uniform probability over a set $\{U\}$. Then, we measure the bit-string $b$ in the computational basis. Taken together, we sample a data point, which is a tuple of the sampled unitary and the measured bit-string $x_{i}=(U_{i},b_{i})$, with probability
@@ E.2 Regularization
In the main text, we numerically explore different regularization strengths. Moreover, we have focused on two measurement schemes: global-$XZ$ measurements, and random-$XZ$ measurements. For the former, we measure bit-strings in two fixed bases $X^{\otimes n}$ and $Z^{\otimes n}$, and we can estimate certain physical observables using empirical means. For the latter, using the toolbox of randomized measurements, we apply random unitaries from an ensemble and measure bit-strings in the computational basis. Depending on the choice of the unitary architectures, different physical properties can be estimated. Performing classical shadow tomography using random-$XZ$ measurements—as detailed in Appendix C—allows us to estimate any observables in the visible space of our measurement. We now discuss our choices of regularizers for the two systems considered in the main text:
@@
$\hat{\rho}_{N}^{B}=\otimes_{i\in B}\left(2U^{\alpha_{i}\dagger}\rvert b_{i}\rangle\!\langle b_{i}\lvert U^{\alpha_{i}}-\frac{\mathbb{I}}{2}\right).$ (118)
@@
$\tilde{\rho}^{B}(A)=\mathop{\mathrm{Tr}}_{\bar{B}}(\rvert\psi(A)\rangle\!\langle\psi(A)\lvert).$ (119)

### v4

@@ B.6 Main result: A concentration inequality for infidelity
For simplicity, we assume that $|\hat{\psi}^{(N)}\rangle$ and $|\phi\rangle$ are normalized elements of the equivalence classes $\hat{\psi}^{(N)}]$ and $\phi]$, respectively. Let $\tilde{g}$ be the Fubini-Study metric of the projective MPS manifold $\tilde{M}$ induced by the Fubini-Study metric of the projective Hilbert space via Eq. (41). On this manifold, $\gamma$ is the shortest connecting geodesic between $\gamma(0)=[\phi]$ and $\gamma(1)=[\hat{\psi}^{(N)}]$. By Proposition. 1, the estimator converges to the target state in the infinite-sample limit
$\hat{\psi}^{(N)}]\overset{\mathrm{p.}}{\longrightarrow}[\phi],$ (64)
@@
${\mbox{$\langle\dot{\gamma}|\dot{\gamma}\rangle$}_{\tilde{g}}}={\sum_{a,b}{\Delta}^{(N)}_{a}\,\tilde{g}_{a,b}\,{\Delta}^{(N)}_{b}}$. (67)
@@
${\mbox{$\langle\dot{\gamma}|\dot{\gamma}\rangle$}_{\tilde{g}}}={{{\Delta}^{(N)}}^{\top}\tilde{g}\,{\Delta}^{(N)}},$ (68)
@@
The $\eta$-term in the denominator is a consequence of the fact that the estimator $\hat{\psi}^{(N)}]$ is only asymptotically normal. The smaller one wants to make $\eta$, the larger $N_{0}$ becomes so that we have to move “closer” to the asymptotic regime. Asymptotically, i.e., in the limit $N\rightarrow\infty$, we can choose $\eta\rightarrow 0$ to arrive at our main result in Theorem 2
@@
Supplementary Figure 1: Infidelity scaling for interpolation between a random and GHZ 33-qubit state. The numerical results illustrate the polynomial scaling of the infidelity with the number of samples NN across a family of states, parametrized by $x$. Here, $x$ ranges from 0 (darkest color, representing the GHZ state) to 1 (lightest color, representing the random state). Dashed lines indicate polynomial fits of the form $1-F=c\times N^{-\alpha}$, with values for $c$ and $\alpha$ provided in the legend.
@@ B.7 Discussions and interpretations
In this subsection, we discuss Theorem 2 and its connection to the numerical results presented in the main text. This theorem demonstrates that the MPS solution of maximum likelihood estimate (MLE) approaches the target state when the number of samples $N$ is sufficiently large, with the infidelity 1F1-F converging at a rate proportional to $1/\sqrt{N}$. This provides a theoretical foundation for understanding the scaling behavior observed in our numerical experiments.
First, we comment on how this theoretical result relates to our numerics in practice. In the derivation of our theorem, we have assumed that our variational space forms a projective MPS manifold. However, in our numerical simulations, we fix the bond dimension across sites, including variational states that are not always full rank, although they do include a projective MPS manifold as a subspace. In addition, several assumptions underpinning our result could be compromised: 1) the Hessian of the negative log-likelihood (NLL) loss function may not behave as expected, or 2) the unitary ensemble is not tomographically complete in our variational space. Notably, while Theorem 1 confirms that our random-$XZ$ measurements are tomographically complete for real and pure states, our variational approach involves complex numbers for enhanced stability in optimization, leading to a space of complex pure states. To overcome this discrepancy, incorporating random-Pauli measurements, which are known to be tomographically complete, could provide a solution. Alternatively, our analysis of the complex projective MPS manifold could be adapted to the setting of a real projective MPS, and one can verify that the necessary conditions on the manifold structure are satisfied in the subspace of real states.
@@
$\mbox{$|\phi\rangle$}=\sqrt{1-x}\mbox{$|\mathrm{GHZ}\rangle$}+\sqrt{x}\mbox{$|\psi_{\chi=2}\rangle$}.$ (86)
@@
Random-XZXZ classical shadows involve preparing some (unknown) state, randomly measuring some $n$-qubit Pauli string composed of only $X$ and $Z$ single-site Paulis, and then post-processing to predict properties of interest. We assume that the chosen Pauli is uniformly sampled from the set of all $n$-qubit strings composed of only $X$s and $Z$s. As this set of measurements is not informationally complete, random-$XZ$ classical shadows do not allow us to predict any arbitrary observable. However, we can predict some observables of interest, namely, those that contribute to our regularization step. In this appendix, we will discuss which observables can be estimated with this technique, discuss how to post-process the data to estimate observables, and how to calculate the sample complexity required for learning Pauli strings. We specifically discuss the sample complexity for learning Pauli strings because these are the observables that we estimate in our regularization step. In the first section of this appendix, we introduce these ideas pedagogically, with the rigorous backings for all our claims provided in the second section.
@@ C.1 Performing random-$XZ$ classical shadow tomography
$U=\otimes_{i=1}^{n}U^{a_{i}}\hskip 2.84526pt\text{ where }a_{i}\in\{X,Z\}.$ (87)
Here, $U^{a_{i}}$ is the single-qubit unitary which rotates from the computational basis to the (different) basis of the Pauli operator $\sigma_{i}^{a}$ acting on site $i$. Therefore, acting on the prepared state, the ensemble $\mathcal{U}_{XZ}$ allows one to randomly measure any $n$-qubit Pauli string composed of only $X$ and $Z$ single-site Paulis. Again, this set of measurements is not informationally complete, and as such, we cannot predict any arbitrary observable with random-$XZ$ classical shadows.
@@
We can estimate observables $O\in\textnormal{{VisibleSpace}}(\mathcal{U}_{XZ})$ by post-processing the data to construct an estimate $\hat{\rho}$ of the true density matrix $\rho$. This estimate is called the “classical shadow” of the true density matrix, and we can use it to compute $\mathop{\mathrm{Tr}}(\hat{\rho}O)\approx\mathop{\mathrm{Tr}}(\rho O)$. We construct the classical shadow using the data from our measurements. Each measurement snapshot can be characterized by the unitary applied, $U=\otimes_{i}\hskip 2.84526ptU^{a_{i}}$, and the computational basis state measured, $\mbox{$|b\rangle$}=\otimes_{i}\mbox{$|b_{i}\rangle$}$. Here the $i$ indexes the qubits. We can construct a single-shot estimate of our density matrix, $\hat{\rho}_{1}$, from random-$XZ$ data as
@@
$=$ $\otimes_{i=1}^{n}\left(\mathop{\mathrm{Tr}}(U^{a_{i}\dagger}\rvert b_{i}\rangle\!\langle b_{i}\lvert U^{a_{i}}X)X+\mathop{\mathrm{Tr}}(U^{a_{i}\dagger}\rvert b_{i}\rangle\!\langle b_{i}\lvert U^{a_{i}}Z)Z+\frac{1}{2}\mathbb{I}\right)$
$=$ $\otimes_{i=1}^{n}\left(2U^{a_{i}\dagger}\rvert b_{i}\rangle\!\langle b_{i}\lvert U^{a_{i}}-\frac{\mathbb{I}}{2}\right).$
@@
Here, we represent the unitary applied in the $j$th random measurement and the outcome bitstring of that measurement as $U^{(j)}$ and $b^{(j)}$, respectively.
@@ C.2 Derivation of tools for random-$XZ$ classical shadows
In the second line, we take the trace over subsystem “BB”, which we define to be the second subsystem (i.e., the one with $P$ in the term $\mathds{1}\otimes P$). In the third line, $\text{\rm Pauli}_{\mathds{1},Z}(n)$ is the set of all normalized Pauli strings containing only $\mathds{1}$s and $Z$s. One can begin to see the transition from line two to line three by considering the single-qubit measurement channel. For a single qubit, we have $\sum_{b\in\{0,1\}}\lvert b,b\rangle\!\langle b,b\rvert=\frac{1}{2}(\mathds{1}\mathds{1}+ZZ)$. This can be extended to $n$ qubits, where the $n$-qubit version $\sum_{b\in\{0,1\}^{n}}\mbox{$|b,b\rangle$}\mbox{$\langle b,b|$}$ becomes a sum of all length-$n$ $\mathds{1}$,$Z$ strings tensored twice, normalized by a factor of $1/2^{n}$ [15]. This normalization does not appear in the final expression above because we absorb it into the two $\tilde{P}$s such that they are all normalized with respect to the Hilbert-Schmidt inner product.
@@
$\mathcal{M}_{XZ}(A)=\left(\mathcal{M}_{XZ}^{(1)}\right)^{\otimes n}(A),\hskip 14.22636pt\textnormal{ where }\hskip 2.84526pt\mathcal{M}_{XZ}^{(1)}(B)=\frac{1}{4}\left(\mathop{\mathrm{Tr}}(BX)X+\mathop{\mathrm{Tr}}(BZ)Z\right)+\frac{1}{2}\mathop{\mathrm{Tr}}(B)\mathbb{I}\hskip 8.53581pt$ (97)
@@
In this appendix, we will explore how and when we can fully characterize real pure states. These states can be expressed as a pure-state vector $\mbox{$|\psi\rangle$}=\sum_{b}c_{b}\mbox{$|b\rangle$}$ with real coefficients $c_{b}^{*}=c_{b}$. For simplicity, we will assume that the basis $\{\mbox{$|b\rangle$}\}_{b\in\{0,1\}^{n}}$ is the computational basis, unless indicated otherwise. We will demonstrate that we can fully characterize these states by only measuring each qubit in either the $X$ or the $Z$ basis (Lemma 5). While our numerical demonstrations are not always limited to measurements only in the $X$ or $Z$ bases, our motivation for specifically pursuing measurement schemes of this type is that randomized-$XZ$ measurements are tomographically complete on the subspace of real and pure states (Theorem 1).
@@ D.2 Proof of Theorem 1
In this subsection, we prove that $XZ$ measurements, where each qubit is measured in either the $X$ or the $Z$ basis, are tomographically complete on the space of real pure states. Tomographic completeness is similar to informational completeness but defined with respect to state characterization. A measurement primitive, which samples $U\sim\mathcal{U}$ and subsequently measures the rotated state in the computational basis $\{\mbox{$|b\rangle$}:b\in\{0,1\}^{n}\}$, is tomographically complete if and only if for each $\rho\neq\sigma$, there exist $U\in\mathcal{U}$ and $b$ such that $\mbox{$\langle b|$}U\rho U^{\dagger}\mbox{$|b\rangle$}\neq\mbox{$\langle b|$}U\sigma U^{\dagger}\mbox{$|b\rangle$}$ [25]. For our purposes, we are only interested in establishing tomographic completeness on a subspace of the density matrices, namely, all states that are both real and pure.
@@
Up to a global sign, we can uniquely characterize any real state vector $|\psi\rangle$ using XZ measurements, formally defined as measuring any qubit in either the X or the Z basis. We represent our real pure state in the computational basis $\{\mbox{$|b\rangle$}\}_{b\in\{0,1\}^{n}}$ as $\mbox{$|\psi\rangle$}=\sum_{b}c_{b}\mbox{$|b\rangle$}$, where $c_{b}^{*}=c_{b}$ and $\sum_{b}|c_{b}|^{2}=1$. We uniquely characterize the state by learning each $c_{b}$ up to a global sign.
@@
Inductive step ($k$ qubits). Assume we can learn all the relative signs of a real, pure state on k1k-1 qubits with $XZ$ measurements. We have a $k$-qubit state. If we measure our $k$th qubit in the computational $Z$ basis, then, by the inductive assumption, we can learn all the relative signs in the $\{\mbox{$|0\rangle$}\otimes\mbox{$|\tilde{b}\rangle$}\}_{\tilde{b}\in\{0,1\}^{k-1}}$ subspace as well as in the $\{\mbox{$|1\rangle$}\otimes\mbox{$|\tilde{b}\rangle$}\}_{\tilde{b}\in\{0,1\}^{k-1}}$ subspace. Here, the tilde on $\tilde{b}$ indicates that the bitstring is on k1k-1 qubits instead of $k$. Also note that when we project the $k$th qubit of our state vector $|\psi\rangle$ onto $|0\rangle$ or $|1\rangle$, the state on the remaining k1k-1 qubits is projected onto a real pure state.
What is, therefore, left to be shown is that we can learn the relative sign between these two subspaces. If the amplitudes of all states in either of the subspaces are zero, then the relative sign is inconsequential because we already know the state up to a global phase. Otherwise, there must exist at least one state in each subspace that has nonzero amplitude. Let us choose two computational basis states: $b^{(0)}$ from the $\{\mbox{$|0\rangle$}\otimes\mbox{$|\tilde{b}\rangle$}\}_{\tilde{b}}$ subspace and $b^{(1)}$ from the $\{\mbox{$|1\rangle$}\otimes\mbox{$|\tilde{b}\rangle$}\}_{\tilde{b}}$ subspace, respectively. We choose them such that they that have nonzero amplitude and, among all possible choices, they have the smallest Hamming weight $\mathbf{d}(b^{(0)},b^{(1)})$; this Hamming weight must be greater than or equal to 1 because the $k$th bit will always differ. We can extract the relative sign of the two spaces by estimating the observable $O\in\textnormal{{VisibleSpace}}(\mathcal{U}_{XZ})$ given by
@@
where $b^{(0)}_{i}$ is the $i$th bit of bitstring $b^{(0)}$, and likewise $b^{(1)}_{i}$ is the $i$th bit of $b^{(1)}$.
@@
A measurement primitive which samples $U\sim\mathcal{U}_{XZ}$ and subsequently measures the rotated state in the computational basis $\{\mbox{$|b\rangle$}:b\in\{0,1\}^{n}\}$ is tomographically complete [25] on the space of real, pure states. Namely, for all states $\lvert\psi\rangle\!\langle\psi\rvert\neq\lvert\phi\rangle\!\langle\phi\rvert$, there exist $U\in\mathcal{U}_{XZ}$ and $b$ such that $\mbox{$\langle b|$}U\rho U^{\dagger}\mbox{$|b\rangle$}\neq\mbox{$\langle b|$}U\sigma U^{\dagger}\mbox{$|b\rangle$}$.
@@ E.1 Maximum likelihood estimation: Randomized Kullback–Leibler divergence
More specifically, for our purposes, the dataset consists of bitstring measurements of a target state $\rvert\phi\rangle\!\langle\phi\lvert$ in the computational basis after applying a unitary $U$. In a randomized measurement scheme, such as the $XZ$-shadow measurements discussed in Appendix C, we first draw a unitary from an ensemble $\mathcal{U}$ with a uniform probability over a set $\{U\}$. Then, we measure the bitstring $b$ in the computational basis. Taken together, we sample a data point, which is a tuple of the sampled unitary and the measured bitstring $x_{i}=(U_{i},b_{i})$, with probability
@@ E.2 Regularization
In the main text, we numerically explore different regularization strengths. Moreover, we have focused on two measurement schemes: global-$XZ$ measurements, and random-$XZ$ measurements. For the former, we measure bitstrings in two fixed bases $X^{\otimes n}$ and $Z^{\otimes n}$, and we can estimate certain physical observables using empirical means. For the latter, using the toolbox of randomized measurements, we apply random unitaries from an ensemble and measure bitstrings in the computational basis. Depending on the choice of the unitary architectures, different physical properties can be estimated. Performing classical shadow tomography using random-$XZ$ measurements—as detailed in Appendix C—allows us to estimate any observables in the visible space of our measurement. We now discuss our choices of regularizers for the two systems considered in the main text:
@@
$\hat{\rho}_{N}^{B}=\otimes_{i\in B}\left(2U^{a_{i}\dagger}\rvert b_{i}\rangle\!\langle b_{i}\lvert U^{a_{i}}-\frac{\mathbb{I}}{2}\right).$ (118)
@@
[…]